# Cones of Noether–Lefschetz divisors and moduli spaces of hyperkähler manifolds

### v2

@@ 4.2. Moduli of ${\rm{Kum}}_{n}$-type hyperkähler manifolds
where $\chi:\widehat{\rm{O}}^{+}\left(\Lambda\right)\longrightarrow\{\pm 1\}$ is the character defined by the action of $\widehat{\rm{O}}^{+}\left(\Lambda\right)$ on $D(\Lambda)$. Let $h=c_{1}(L)\in\Lambda$, with $\langle h,h\rangle=2d$ and divisibility $\gamma$. Since $\widetilde{\rm{SO}}^{+}\left(\Lambda\right)\subset{\rm{Mon}}^{2}\left(\Lambda\right)$, up to monodromy one can always assume $h=\gamma(e+tf)-a\delta$ for appropriate $t$ and $a$, where $\delta$ is the generator of $A_{1}(-(n+1))$.
For $\gamma=1,2$, the lattice $\Lambda_{h}$ is in the form $U^{\oplus 2}\oplus Q_{d}$ (resp. $Q_{t}$ with $d=4t-(n+1)$) where
@@
${\rm{Mon}}^{2}\left(\Lambda,h\right)=\langle\widetilde{\rm{SO}}^{+}(\Lambda_{h}),\sigma_{\kappa}\rangle,\;\;\hbox{ where }\;\;\kappa=\begin{cases}\delta&\mbox{ if }\gamma=1,\\ v&\mbox{ if }\gamma=2.\\ \end{cases}$
@@
Observe that, for any $g\in O\left(\Lambda,h\right)$, we have $\det(g)=\det(g|_{\Lambda_{h}})$. Then the statement is essentially [ABL24]*Lemma 3.7. For $\gamma=1,2$ we also need to prove that $\sigma_{\kappa}\in{\rm{Mon}}^{2}\left(\Lambda\right)$ via the restriction: since $\det(\sigma_{\kappa})=-1$, this is equivalent to prove that, if we see $\kappa$ as an element of $\Lambda$, we have $-\sigma_{\kappa}\in\widetilde{\rm{O}}^{+}(\Lambda)$ i.e. $\chi(\sigma_{\kappa})=-1$. Since $\kappa=3(\gamma-1)f-\delta$, this can be checked via an explicit computation. ∎
For $\gamma=1,2$, the period domain $\mathcal{M}^{\gamma}_{{\rm{Kum}}_{n},2d}=\mathcal{D}_{\Lambda_{h}}/{\rm{Mon}}^{2}\left(\Lambda,h\right)$ is a partial compactification of the moduli space of hyperkähler $2n$-folds of generalized Kummer type with a primitive polarization of degree $2d$ and divisibility $\gamma$. It is always irreducible [Ono22] and never empty for $\gamma=1$ (the split case). When $\gamma=2$ it is non-empty only for $d\equiv-(n+1)\pmod{4}$.
@@
Note that, under our hypothesis, $-\sigma_{\kappa}\in\widetilde{\rm{O}}^{+}\left(\Lambda_{h}\right)$. For $\gamma=1$, this holds since $\ell_{*}=-\ell_{*}$. For $\gamma=2$, observe that $|D(\Lambda_{h})|=d\cdot(n+1)=n+1$, hence $D(\Lambda_{h})=\langle\kappa_{*}\rangle$ since $\kappa=v$ is primitive with divisibility $n+1$; clearly $\sigma_{\kappa}(\kappa_{*})=-\kappa_{*}$. Now we prove (22) under the more general hypothesis that $-\sigma_{\kappa}\in\widetilde{\rm{O}}^{+}\left(\Lambda_{2}\right)$ i.e. $\chi(\sigma_{\kappa})=-1$.
We can write $\widehat{\rm{O}}^{+}\left(\Lambda_{h}\right)=\bigcup_{i,j\in\{-1,+1\}}M_{i,j}$, where $M_{i,j}$ is the set of isometries $g\in\widehat{\rm{O}}^{+}\left(\Lambda_{2}\right)$ such that $(\chi(g),\det(g))=(i,j)$. Clearly $\widetilde{\rm{SO}}^{+}(\Lambda_{h})=M_{+1,+1}$ and, under our hypothesis, $\sigma_{\kappa}\cdot\widehat{\rm{SO}}^{+}(\Lambda_{h})=M_{-1,-1}$. By Lemma 4.6 then ${\rm{Mon}}^{2}\left(\Lambda,h\right)=M_{+1,+1}\cup M_{-1,-1}$. Now $-\operatorname{Id}\in M_{-1,1}$, since $\Lambda_{h}$ has even rank, hence $-\operatorname{Id}\cdot M_{i,j}=M_{-i,j}$ and (22) follows. ∎
@@
    when $\gamma=2$ for $t\leq 11$ and $t=13,15,17,19$, where $n=4t-2$.
@@
As in the proof of Theorem 4.5, we show that $X=\mathcal{D}_{\Lambda_{h}}\big{/}\widetilde{\rm{O}}^{+}(\Lambda_{h})$ is uniruled and by Lemma 4.7 we conclude uniruledness for $\mathcal{M}_{{\rm{Kum}}_{2},2}^{\gamma}$. By Proposition 1.2, we just need to verify that $04c_{0,0}(E_{3,\Lambda_{h}})+\frac{1}{4}c_{-1,0}(E_{3,\Lambda_{h}})<$ 0. Again, we exhibit only one case. If $E_{3,\Lambda_{h}}=\sum_{\mu\in D(\Lambda_{h})}E_{\mu}(q)e_{\mu}$, one computes [weilrep]:
$E_{0}(q)=\begin{cases}1-\frac{4250}{263}q-\frac{12600}{263}q^{2}+O(q^{3})&\hbox{ if $n=20$ and $\gamma=1$}\\ 1-\frac{130}{7}q-\frac{288}{7}q^{2}+O(q^{3})&\hbox{ if $n=4t-2$ with $t=19$ and $\gamma=2$}.\end{cases}$
@@
$\mathcal{M}^{2}_{{\rm{Kum}}_{2},2}=\mathcal{D}_{\Lambda_{h}}\big{/}\widetilde{\rm{O}}^{+}\left(\Lambda_{h}\right),$
where $\Lambda_{h}=U^{\oplus 2}\oplus A_{2}(-1)$, is known to be rational [WW21]*Theorem 5.4. More concretely, there is a finite union of Heegner divisors $\mathcal{H}$, see [WW21]*Equation 5.8, such that the algebra of meromorphic modular forms ${\rm{M}}_{*}^{!}\left(\widetilde{\rm{O}}^{+}(\Lambda_{h}),\mathcal{H}\right)$, that is, meromorphic sections of $\lambda^{\otimes k}$ with $k\in\mathbb{Z}$ and poles supported along $\mathcal{H}$ is finitely generated by forms of positive weight. By work of Looijenga [Loo03] the projective variety $\widehat{X}={\rm{Proj}}\left(\oplus_{k\geq 0}{\rm{M}}_{k}^{!}\left(\widetilde{\rm{O}}^{+},\mathcal{H}\right)\right)$ is a compactification of $\mathcal{D}\big{/}\widetilde{\rm{O}}^{+}-\mathcal{H}$ that interpolates between the Baily–Borel and toroidal compactifications. When the generators are relation-free, as it is shown in [WW21] for $\Lambda_{h}=U^{\oplus 2}\oplus A_{2}(-1)$, the resulting ring is a polynomial algebra with generators of mixed weights. In this case $\widehat{X}$ is a weighted projective space, in particular rational. The same holds for some of the first OG6 cases. Indeed if $\Lambda_{h}=U^{\oplus 2}\oplus A_{1}(-1)^{\oplus 3}$ or $\Lambda_{h}=U^{\oplus 2}\oplus A_{1}(-1)^{\oplus 1}\oplus A_{2}(-1)$, then [WW21]*Theorem 5.4 implies that the resulting modular varieties $\mathcal{D}_{\Lambda_{h}}\big{/}\widetilde{\rm{O}}^{+}(\Lambda_{h})$ are also rational. We summarize the results relevant for this paper:
@@
For $\gamma=3,6$, when non-empty, the canonical class of every component $\mathcal{M}$ of the moduli space $\mathcal{M}_{{\rm{Kum}}_{2},2d}^{\gamma}$ is given by
@@
In particular, it lies in the interior of the NL cone and it has positive intersection with any complete curve not intersecting the boundary of a toroidal compactification.
@@
By Lemma 4.6, the branch divisor of the modular projection $\pi:\mathcal{D}_{\Lambda_{h}}\longrightarrow\mathcal{D}_{\Lambda_{h}}\big{/}{\rm{Mon}}^{2}\left(\Lambda,h\right)$ is trivial, since both $\sigma_{\rho}$ and $-\sigma_{\rho}$ have negative determinant on a lattice of even rank, see (14) and its surrounding discussion. ∎
@@
In the case of the rational moduli space $\left(\mathcal{M}_{{\rm{Kum}}_{2},2}^{2}\right)^{\circ}$ parameterizing polarized hyperkähler fourfolds with polarization of degree $2$ and divisibility $2$ the rational Picard group ${\rm Pic}_{\mathbb{Q}}\left(\mathcal{M}_{{\rm{Kum}}_{2},2}^{2}\right)$ is one-dimensional, since the space of cusp forms $S_{3,\Lambda_{h}}$ is trivial (this can be computed using [weilrep]). In particular, in this case we have the equality
@@
(23) $\left(\mathcal{M}_{{\rm{Kum}}_{2},2}^{2}\right)^{\circ}\longrightarrow\mathcal{D}_{\Lambda_{h}}\big{/}{\rm{Mon}}^{2}\left(\Lambda,h\right)=\mathcal{D}_{\Lambda_{h}}\big{/}\widetilde{\rm{O}}^{+}\left(\Lambda_{h}\right)=\mathcal{M}_{{\rm{Kum}}_{2},2}^{2}.$
Since ${\rm Pic}_{\mathbb{Q}}\left(\mathcal{M}_{{\rm{Kum}}_{2},2}^{2}\right)$ is one-dimensional, it is thus enough to show that the complement of $\left(\mathcal{M}_{{\rm{Kum}}_{2},2}^{2}\right)^{\circ}$ in $\mathcal{M}_{{\rm{Kum}}_{2},2}^{2}$ contains a primitive Heegner divisor $P_{\rho}$ since then $P_{\rho}$ is a positive rational multiple of the Hodge class $\lambda$ and so is ample, meaning that $\left(\mathcal{M}_{{\rm{Kum}}_{2},2}^{2}\right)^{\circ}$ is an open subset of the complement of a hyperplane in the Baily-Borel compactification $\mathcal{D}_{\Lambda_{h}}\big{/}\widetilde{\rm{O}}^{+}\left(\Lambda_{h}\right)^{\rm BB}\subset\mathbb{P}^{N}$. This exactly means that $\left(\mathcal{M}_{{\rm{Kum}}_{2},2}^{2}\right)^{\circ}$ is quasi-affine.
@@
with $\Lambda=U^{\oplus 3}\oplus A_{1}(-3)$. We call $\delta$ the generator of the last factor and $h=c_{1}(H)$. By [Yos16], see also [MTW2018]*Page 452, an ample class $h$ cannot lie in the orthogonal complement in $H^{1,1}(X,\mathbb{R})$ of classes $\rho\in{\rm{NS}}(X)$ whose square is $-6$ and divisibility in $H^{2}(X,\mathbb{Z})$ is $2,3$ or $6$. In particular, if such a class is orthogonal to $h$, then $D_{\rho}$ defines a hyperplane in $\mathcal{D}_{\Lambda_{h}}$ and the image of the period map misses the corresponding divisor $P_{\rho}$. Singling out classes in $H^{1,1}(X,\mathbb{Z})$ whose orthogonal complements give the chamber decomposition of the positive cone $C(X)\subset H^{1,1}(X,\mathbb{R})$ is a general method to describe the complement of the image of the period map, see for instance [DM19]*Theorem 6.1. Thus, it is enough to show that there exists an integral class $\rho\in\Lambda_{h}$ of square $\langle\rho,\rho\rangle=-6$ and divisibility in $\Lambda$ given by ${\rm{div}}_{\Lambda}(\rho)\in\{2,3,6\}$. Since $\widetilde{\rm{SO}}^{+}(\Lambda)\subset{\rm{Mon}}^{2}(\Lambda)$, one can assume $h=2(e+f)-\delta$, and taking $\rho=3f-\delta$ one has the desired property. In this case the divisibility in $\Lambda$ is $3$, and the missed primitive Heegner divisor in $\mathcal{M}_{{\rm{Kum}}_{2},2}^{2}$ is $P_{\rho}=P_{-\frac{1}{3},\frac{v}{3}}$. ∎
@@
We have seen that one can compute the NL cone of $\mathcal{M}_{{\rm{Kum}}_{2},2}^{2}$ just from the fact that $\dim_{\mathbb{Q}}{\rm Pic}_{\mathbb{Q}}\left(\mathcal{M}_{{\rm{Kum}}_{2},2}^{2}\right)=1$. The result of Lemma 4.6 a priori means that one can also compute the NL cone of the moduli space $\mathcal{M}_{{\rm{Kum}}_{2},2}^{1}$ using the method of Section 3 for the quotient $\mathcal{D}_{\Lambda_{h}}\big{/}\widetilde{\rm{O}}^{+}\left(\Lambda_{h}\right)$. In this case, however, the bound obtained in Theorem 3.7 is too large for this to be computationally feasible. A computation of the cone generated by all $P_{\Delta,\delta}$ with $\Delta\leq 10000$ yields the conjectural description

### v3

@@ 4.2. Moduli of ${\rm{Kum}}_{n}$-type hyperkähler manifolds
where $\widehat{\rm{O}}^{+}\left(\Lambda\right)$ is the group of orientation-preserving isomorphisms of $\Lambda$ which act as $\pm{\rm Id}$ on $D(\Lambda)$ and $\chi:\widehat{\rm{O}}^{+}\left(\Lambda\right)\longrightarrow\{\pm 1\}$ is the character defined by the action of $\widehat{\rm{O}}^{+}\left(\Lambda\right)$ on $D(\Lambda)$. Let $h=c_{1}(L)\in\Lambda$, with $\langle h,h\rangle=2d$ and divisibility $\gamma$. Since $\widetilde{\rm{SO}}^{+}\left(\Lambda\right)\subset{\rm{Mon}}^{2}\left(\Lambda\right)$, up to monodromy one can always assume $h=\gamma(e+tf)-a\delta$ for appropriate $t$ and $a$, where $\delta$ is the generator of $A_{1}(-(n+1))$.
For $\gamma=1,2$, the lattice $\Lambda_{h}$ is in the form $U^{\oplus 2}\oplus Q_{d}$ (resp. $Q_{t}$ with d=4t(n+1)d=4t-(n+1)) where
@@
[…]
@@
Observe that, for any $g\in{\rm O}\left(\Lambda,h\right)$, we have $\det(g)=\det(g|_{\Lambda_{h}})$. Then the statement is essentially [ABL24]*Lemma 3.7. For $\gamma=1,2$ we also need to prove that $\sigma_{\kappa}\in{\rm{Mon}}^{2}\left(\Lambda\right)$ via the restriction: since $\det(\sigma_{\kappa})=-1$, this is equivalent to proving that, if we see $\kappa$ as an element of $\Lambda$, we have $\sigma_{\kappa}\in\widetilde{\rm{O}}^{+}(\Lambda)$ i.e. $\chi(\sigma_{\kappa})=-1$. Since $\kappa=3(\gamma-1)f-\delta$, this can be checked via an explicit computation. ∎
For $\gamma=1,2$, the period domain $\mathcal{M}^{\gamma}_{{\rm{Kum}}_{n},2d}=\mathcal{D}/{\rm{Mon}}^{2}\left(\Lambda,h\right)$ is a partial compactification of the moduli space of hyperkähler $2n$-folds of generalized Kummer type with a primitive polarization of degree $2d$ and divisibility $\gamma$. It is always irreducible [Ono22] and never empty for $\gamma=1$ (the split case). When $\gamma=2$ it is non-empty only for $d\equiv-(n+1)\pmod{4}$.
@@
Note that, under our hypothesis, $\sigma_{\kappa}\in\widetilde{\rm{O}}^{+}\left(\Lambda_{h}\right)$. For $\gamma=1$, this holds since $\ell_{*}=$-$\ell_{*}$. For $\gamma=2$, observe that $|D(\Lambda_{h})|=d\cdot(n+1)=n+1$, hence $D(\Lambda_{h})=\langle\kappa_{*}\rangle$ since $\kappa=v$ is primitive with divisibility $n+1$; clearly $\sigma_{\kappa}(\kappa_{*})=-\kappa_{*}$. Now we prove (22) under the more general hypothesis that $\sigma_{\kappa}\in\widetilde{\rm{O}}^{+}\left(\Lambda_{2}\right)$ i.e. $\chi(\sigma_{\kappa})=-1$.
We can write $\widehat{\rm{O}}^{+}\left(\Lambda_{h}\right)=\bigcup_{i,j\in\{-1,+1\}}M_{i,j}$, where $M_{i,j}$ is the set of isometries $g\in\widehat{\rm{O}}^{+}\left(\Lambda_{2}\right)$ such that $(\chi(g),\det(g))=(i,j)$. Clearly $\widetilde{\rm{SO}}^{+}(\Lambda_{h})=M_{+1,+1}$ and, under our hypothesis, $\sigma_{\kappa}\cdot\widehat{\rm{SO}}^{+}(\Lambda_{h})=M_{-1,-1}$. By Lemma 4.6 then ${\rm{Mon}}^{2}\left(\Lambda,h\right)=M_{+1,+1}\cup M_{-1,-1}$. Now $\operatorname{Id}\in M_{-1,1}$, since $\Lambda_{h}$ has even rank, hence $\operatorname{Id}\cdot M_{i,j}=M_{-i,j}$ and (22) follows. ∎
@@
    when $\gamma=2$ for $t\leq 11$ and $t=13,15,17,19$, where n=4t2n=4t-2.
@@
As in the proof of Theorem 4.5, we show that $X=\mathcal{D}\big/\widetilde{\rm{O}}^{+}(\Lambda_{h})$ is uniruled and by Lemma 4.7 we conclude uniruledness for $\mathcal{M}_{{\rm{Kum}}_{2},2}^{\gamma}$. By Proposition 1.2, we just need to verify that $4c_{0,0}(E_{3,\Lambda_{h}})+\frac{1}{4}c_{-1,0}(E_{3,\Lambda_{h}})<0$. Again, we exhibit only one case. If $E_{3,\Lambda_{h}}=\sum_{\mu\in D(\Lambda_{h})}E_{\mu}(q)e_{\mu}$, one computes [weilrep]:
[…]
@@
$\mathcal{M}^{2}_{{\rm{Kum}}_{2},2}=\mathcal{D}\big/\widetilde{\rm{O}}^{+}\left(\Lambda_{h}\right),$
[…]
@@
If $\gamma=3,6$ and the moduli space $\mathcal{M}_{{\rm{Kum}}_{2},2d}^{\gamma}$ is non-empty, then the canonical class of any component $\mathcal{M}$ of the moduli space $\mathcal{M}_{{\rm{Kum}}_{2},2d}^{\gamma}$ is given by
@@
In particular, $K_{\mathcal{M}}$ lies in the interior of the NL cone and it has positive intersection with any complete curve not intersecting the boundary of a toroidal compactification.
@@
By Lemma 4.6, the branch divisor of the modular projection $\pi:\mathcal{D}\longrightarrow\mathcal{D}\big/{\rm{Mon}}^{2}\left(\Lambda,h\right)$ is trivial, since both $\sigma_{\rho}$ and $\sigma_{\rho}$ have negative determinant on a lattice of even rank, see (14) and its surrounding discussion. ∎
@@
In the case of the rational moduli space $\left(\mathcal{M}_{{\rm{Kum}}_{2},2}^{2}\right)^{\circ}$, parameterizing polarized hyperkähler fourfolds with polarization of degree $2$ and divisibility $2$, the rational Picard group ${\rm Pic}_{\mathbb{Q}}\left(\mathcal{M}_{{\rm{Kum}}_{2},2}^{2}\right)$ is one-dimensional, since the space of cusp forms $S_{3,\Lambda_{h}}$ is trivial (this can be computed using [weilrep]). In particular, in this case we have the equality
@@
(23) $\left(\mathcal{M}_{{\rm{Kum}}_{2},2}^{2}\right)^{\circ}\longrightarrow\mathcal{D}\big/{\rm{Mon}}^{2}\left(\Lambda,h\right)=\mathcal{D}\big/\widetilde{\rm{O}}^{+}\left(\Lambda_{h}\right)=\mathcal{M}_{{\rm{Kum}}_{2},2}^{2}.$
Since ${\rm Pic}_{\mathbb{Q}}\left(\mathcal{M}_{{\rm{Kum}}_{2},2}^{2}\right)$ is one-dimensional, it is thus enough to show that the complement of $\left(\mathcal{M}_{{\rm{Kum}}_{2},2}^{2}\right)^{\circ}$ in $\mathcal{M}_{{\rm{Kum}}_{2},2}^{2}$ contains a primitive Heegner divisor $P_{\rho}$ since then $P_{\rho}$ is a positive rational multiple of the Hodge class $\lambda$ and so is ample, meaning that $\left(\mathcal{M}_{{\rm{Kum}}_{2},2}^{2}\right)^{\circ}$ is an open subset of the complement of a hyperplane in the Baily-Borel compactification $\mathcal{D}\big/\widetilde{\rm{O}}^{+}\left(\Lambda_{h}\right)^{\rm BB}\subset\mathbb{P}^{N}$. This exactly means that $\left(\mathcal{M}_{{\rm{Kum}}_{2},2}^{2}\right)^{\circ}$ is quasi-affine.
@@
with $\Lambda=U^{\oplus 3}\oplus A_{1}(-3)$. We call the generator of the last factor $\delta$ and let $h=c_{1}(H)$. By [Yos16], see also [MTW2018]*Page 452, an ample class $h$ cannot lie in the orthogonal complement in $H^{1,1}(X,\mathbb{R})$ of classes $\rho\in{\rm{NS}}(X)$ whose square is 6-6 and divisibility in $H^{2}(X,\mathbb{Z})$ is $2,3$ or $6$. In particular, if such a class is orthogonal to $h$, then $D_{\rho}$ defines a hyperplane in $\mathcal{D}$ and the image of the period map misses the corresponding divisor $P_{\rho}$. Singling out classes in $H^{1,1}(X,\mathbb{Z})$ whose orthogonal complements give the chamber decomposition of the positive cone $C(X)\subset H^{1,1}(X,\mathbb{R})$ is a general method to describe the complement of the image of the period map, see for instance [DM19]*Theorem 6.1. Thus, it is enough to show that there exists an integral class $\rho\in\Lambda_{h}$ of square $\langle\rho,\rho\rangle=-6$ and divisibility $2,3,$ or $6$ in $\Lambda$. Since $\widetilde{\rm{SO}}^{+}(\Lambda)\subset{\rm{Mon}}^{2}(\Lambda)$, one can assume $h=2(e+f)-\delta$, and taking $\rho=3f-\delta$ one has the desired property. In this case the divisibility in $\Lambda$ is $3$, and the missed primitive Heegner divisor in $\mathcal{M}_{{\rm{Kum}}_{2},2}^{2}$ is $P_{\rho}=P_{-\frac{1}{3},\frac{v}{3}}$. ∎
@@
We have seen that one can compute the NL cone of $\mathcal{M}_{{\rm{Kum}}_{2},2}^{2}$ just from the fact that $\dim_{\mathbb{Q}}{\rm Pic}_{\mathbb{Q}}\left(\mathcal{M}_{{\rm{Kum}}_{2},2}^{2}\right)=1$. The result of Lemma 4.6 a priori means that one can also compute the NL cone of the moduli space $\mathcal{M}_{{\rm{Kum}}_{2},2}^{1}$ using the method of Section 3 for the quotient $\mathcal{D}\big/\widetilde{\rm{O}}^{+}\left(\Lambda_{h}\right)$. In this case, however, the bound obtained in Theorem 3.7 is too large for this to be computationally feasible. A computation of the cone generated by all $P_{\Delta,\delta}$ with $\Delta\leq 10000$ yields the conjectural description